OpenStudy (anonymous):

Can somebody help me and explain how to do these?1?

OpenStudy (anonymous):

1. What is the discontinuity of the function \[f(x)=\frac{ x^2-4x-12 }{ x+2 }\] A. (-6, 0) B. (6, 0) C. (-2, -8) D. (2, -4) 3. What is the discontinuity and zero of the function \[f(x)=\frac{ 3x^2+x-4 }{ x-1 }\] A. Discontinuity at (-1, 1), zero at \[(\frac{ 4 }{ 3 },0)\] B. Discontinuity at (-1, 1), zero at \[(-\frac{ 4 }{ 3 },0)\] C. Discontinuity at (1, 7), zero at \[(\frac{ 4 }{ 3 },0)\] D. Discontinuity at (1, 7), zero at \[(-\frac{ 4 }{ 3 },0)\] * I know the zero is at \[(-\frac{ 4 }{ 3 },0)\]

OpenStudy (anonymous):

@phi Can you help me please?

OpenStudy (turingtest):

division by 0 is undefined what value of x results in this expression being divided by 0?

OpenStudy (anonymous):

2